OpenStudy (danise):

Find the real or imaginary solution of each equation by factoring X3 + 64 = 0

jimthompson5910 (jim_thompson5910):

x^3 + 64 = 0 x^3 + 4^3 = 0 (x+4)*(x^2 - x*4 + 4^2) = 0 ... using sum of cubes factoring rule here (x+4)*(x^2 - 4x + 16) = 0 I'll let you finish

OpenStudy (danise):

Dude your amazing